the wavelength of x-rays which gives a diffraction angle  2theta is = 16.80degree for a first order diffraction for crystal with interplanar distance 0.200nm?


 

Grade:

1 Answers

Sunil Kumar FP
9 years ago
the braggs law for scattering is given by
[n\lambda = 2d\sin\theta \,,]

wherenis an integer,λis thewavelengthof incident wave,dis the spacing between the planes in the atomic lattice, andθis the angle between the incident ray and the scattering planes

here theta=16.8/2=8.4 degree
n=1
lambda=to calculate
d=2*10^-10 A
lambda=2*2*10^-10sin8.4/1
=.584A